WORD PROBLEM (AGE PROBLEM)
1. Six years ago Jun was 4 times as old as Joan.In four years, he would be twice as old as Joan. How old are they now?
2. At present, Pedro's age is 25percent of his father's age.Twenty years from now Pedro's age would be 50percent of his fathers age. How old are they?
pls help me to solve this problem.. thank you

Answer by mananth(16949)

6 years ago
joan x
jun 4x
..
after 4 years
joan x+10
Jun 4x+10
4x+10=2(x+10)
4x+10=2x+20
4x-2x=20-10
2x=10
x=5 joan 6 years ago. Now age = 11
Jun 4x, = 20, 6 years ago. Now age =26 years now
.......
Let pedro's age be x ( pedro's age is 1/4 of father'e age OR father is 4 times pedros age)
father's age = 4x
..
20 years later
pedro = x+20
Father = 4x+20
2(x+20)=4x+20
2x+40 =4x+20
2x=20
x=10 pedro's age now
father= 10*4 = 40 years